A rectangular box measures 4 ft. by 4 ft. by 3 ft.
Will you be able to fit a 6 ft pole inside the box?

Answer by unlockmath(1688)

Hello,
The answer is Yes.
To figure it out first find the hypotenuse of the bottom of the box.
4^2+4^2=x^2
x=sq rt 32
Now find the hypotenuse of the diagonal inside the box.
(Sq rt 32)^2 + 3^2 = x^2
32 + 9 = x^2
x= sq rt 41 or rewritten as:
x= 6.4 feet (approx)
Make sense?
RJ
www.math.unlock.com